OpenStudy (anonymous):

Will medal and fan.. please help C/4 - 5 = 4 says i need to show 3 steps

OpenStudy (anonymous):

it only takes two steps a) add 5 b) multiply by 4

OpenStudy (anonymous):

not sure what you mean by add 5? what am i adding 5 to?

OpenStudy (anonymous):

both sides, to get rid of the \(-5\) on the left

OpenStudy (anonymous):

what about the c/4?

OpenStudy (anonymous):

what is step b)?

OpenStudy (anonymous):

so lost on this

OpenStudy (anonymous):

if \[\frac{c}{4}=9\] then what is \(c\)?

OpenStudy (anonymous):

it just says to solve c divided by 4 - 5 = 4

OpenStudy (anonymous):

c is 9 ?

OpenStudy (anonymous):

\[\frac{c}{4}-5=4\] right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

add 5 to both sides, what do you get ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

yeah you get \[\frac{c}{4}=9\]

OpenStudy (anonymous):

what number divided by 4 is 9?

OpenStudy (anonymous):

which is another way of asking what is \(9\times 4\)

OpenStudy (anonymous):

36

OpenStudy (anonymous):

still not understanding how to write these steps

OpenStudy (anonymous):

\[\frac{c}{4}-5=4\\ \frac{c}{4}-5+5=4+5\\ \frac{c}{4}=9\\ 4\times \frac{c}{4}=4\times9\\ c=36\]